# Angular momentum sensitivities in scalar-tensor theories

### v1

@@ I Introduction
Here, we propose a novel way to use pulsar data to test the presence of a possible ultralight gravitational scalar field. In more details, the basic idea, which we will develop below, is that the conformal coupling renormalizes the local value of the Newton constant experienced by a pulsar. This in turn changes the pulsar gravitational binding energy, and therefore the pulsar’s moment of inertia and rotational frequency. Indeed, this change is modulated by the scalar field oscillations far from the pulsar, which lie in the nHz band of pulsar-timing arrays for $m_{\varphi}\sim 10^{-22}$ eV. In this paper, we will develop the calculation of this effect from first principles, and show that it can be parametrized by angular momentum “sensitivies” (or “charges”) akin to those describing the scalar fluxes from compact binaries. In a companion paper, we will apply the idea to data from the European Pulsar Timing Array [40] to constrain the theory.
@@ II.1 Theory
Although moments of inertia for rotating neutron stars have recently been computed in several previous works [44, 45, 46, 47, 48, 49], we are only aware of a few older articles where the angular momentum sensitivities are mentioned [7, 10, 50]. However, these works were focused on FJBD theory: to our knowledge, angular momentum sensitivities were never computed in the DEF case. In this work, we will thus update the computed values of angular momentum sensitivities for both theories using modern equations of state for nuclear matter. We will also clarify the weak field limit of the angular momentum sensitivities, and relate the latter to the EOS in section III.4.
@@ IV.1 Fierz-Jordan-Brans-Dicke theory
Let us stress that while neutron stars close to the maximum mass can be formed by stellar evolution, population studies reveal that neutron stars form with masses $\tilde{M}\gtrsim 1M_{\odot}$, much above the theoretically allowed minimum mass $\tilde{M}_{\mathrm{min}}\sim 0.1-0.2M_{\odot}$ [66, 67]. Thus, while the divergence of $s_{I}$ at the lower mass end is an interesting phenomenon, it is not expected to be observationally relevant.
@@ IV.2 Damour and Esposito-Farèse theory
Quite expectedly, spontaneous scalarization also leads to large angular momentum sensitivities, as can be seen from the upper panel of Fig 2, and from Fig 5, where we show the dependence of $s_{I}$ on $\beta$ as spontaneous scalarization gets activated. Additionally, for the same reason as in FJBD theory, $s_{I}$ diverges at the upper and lower mass ends for the three realistic EOSs considered. The low-mass (Newtonian) limit for the polytropic EOS is finite and gives a value $s_{I}\simeq 1+\beta$ in accordance with Eq. (45). To further illustrate this point, we plot in Fig. 6 the mass and angular momentum sensitivities for a positive value $\beta=6.0$. As can be seen, $s_{I}$ does not vanish in the Newtonian limit, not even for a polytropic EOS (which yields the expected value $s_{I}=7$ at low masses).
@@ V Conclusions
Among the possible avenues to test our theoretical results, a prominent role is played by millisecond pulsars, for which the rotational frequency is measured very accurately. In a companion upcoming paper, we constrain possible variations of the latter with the data of the European Pulsar Timing Array experiment [68]. It remains to be seen if the changes in the moment of inertia that we find may be amplified in the dynamical regime of binary pulsars (c.f. the induced and dynamical scalarization of binaries found in Refs. [42, 69], which generalizes the concept of mass sensitivities to binaries).

### v2

@@ I Introduction
Here, we propose a novel way to use pulsar data to test the presence of a possible ultralight gravitational scalar field. In more details, the basic idea, which we will develop below, is that the conformal coupling renormalizes the local value of the Newton constant experienced by a pulsar. This in turn changes the pulsar gravitational binding energy, and therefore the pulsar’s moment of inertia and rotational frequency. Indeed, this change is modulated by the scalar field oscillations far from the pulsar, which lie in the nHz band of pulsar-timing arrays for $m_{\varphi}\sim 10^{-22}$ eV. In this paper, we will develop the calculation of this effect from first principles, and show that it can be parametrized by angular momentum “sensitivies” (or “charges”) akin to those describing the scalar fluxes from compact binaries. In a companion paper [40], we will apply the idea to data from the European Pulsar Timing Array [41] to constrain the theory.
@@ II.1 Theory
Although moments of inertia for rotating neutron stars have recently been computed in several previous works [45, 46, 47, 48, 49, 50], we are only aware of a few older articles where the angular momentum sensitivities are mentioned [7, 10, 51, 52]. In this work, we will thus update the computed values of angular momentum sensitivities for both theories using modern equations of state for nuclear matter. We will also clarify the weak field limit of the angular momentum sensitivities, and relate the latter to the EOS in section III.4.
@@ IV.1 Fierz-Jordan-Brans-Dicke theory
Let us stress that while neutron stars close to the maximum mass can be formed by stellar evolution, population studies reveal that neutron stars form with masses $\tilde{M}\gtrsim 1M_{\odot}$, much above the theoretically allowed minimum mass $\tilde{M}_{\mathrm{min}}\sim 0.1-0.2M_{\odot}$ [68, 69]. Thus, while the divergence of $s_{I}$ at the lower mass end is an interesting phenomenon, it is not expected to be observationally relevant. Finally, notice that any experimental constraint on the coupling constant $\alpha$ using results from this article will be partly degenerate with changes in the equation of state of nuclear matter, see our companion paper [40] for more details.
@@ IV.2 Damour and Esposito-Farèse theory
Quite expectedly, spontaneous scalarization also leads to large angular momentum sensitivities, as can be seen from the upper panel of Fig 2, and from Fig 5, where we show the dependence of $s_{I}$ on $\beta$ as spontaneous scalarization gets activated. Let us give here a quick order-of-magnitude estimate of the constraints that could be obtained on $\beta$ from Pulsar Timing Array (PTA) observations. When requiring the scalar to be an ultra-light dark matter candidate of mass $m$, its oscillations will produce an additional time residual $\Delta t\sim\beta s_{I}\varphi(t)^{2}/m$. However, we know that a similar residual $\Delta t\sim\varphi(t)^{2}/m$ due to the propagation of the pulsar signals in a perturbed metric has been constrained in [36, 38]. By requiring that our additional time residual is smaller than the one constrained in [36, 38], we find the constraint $|\beta s_{I}|\lesssim 1$ in the relevant mass range. A more quantitative analysis is performed in our companion paper [40], to which we refer for more details.
Additionally, for the same reason as in FJBD theory, $s_{I}$ diverges at the upper and lower mass ends for the three realistic EOSs considered. The low-mass (Newtonian) limit for the polytropic EOS is finite and gives a value $s_{I}\simeq 1+\beta$ in accordance with Eq. (45). To further illustrate this point, we plot in Fig. 6 the mass and angular momentum sensitivities for a positive value $\beta=6.0$. As can be seen, $s_{I}$ does not vanish in the Newtonian limit, not even for a polytropic EOS (which yields the expected value $s_{I}=7$ at low masses).
@@ V Conclusions
Among the possible avenues to test our theoretical results, a prominent role is played by millisecond pulsars, for which the rotational frequency is measured very accurately. In a companion paper [40], we constrain possible variations of the latter with the data of the European Pulsar Timing Array experiment [70]. It remains to be seen if the changes in the moment of inertia that we find may be amplified in the dynamical regime of binary pulsars (c.f. the induced and dynamical scalarization of binaries found in Refs. [43, 71], which generalizes the concept of mass sensitivities to binaries).